A stock S has value S = 2 today and in one year the value either doubles or halves. Assuming interest rates are zero: (a) Show how to replicate a call option with strike at 2.5 using the stock and the riskless asset and calculate the price of the option today (b) Similarly replicate the put option struck at 2.5 and calculate its value today.
Answer #1

Current stock price, S0 = 2;

Stock price in up state, Su = 2 x S0 = 2 x 2 = 4

Stock price in down state, Sd = (1/2) x S0 = (1/2) x 2 = 1

K, strike price = 2.5

Part (a); Replicating portfolio for the call option

let's create the replicating portfolio as:

  1. Borrow an amount B at risk free rate, r = 0 (question states assume interest rates are zero)
  2. Buy N number of underlying stock

Value of the replicating portfolio at the end of 1 year in the up state:

  1. Borrowed amount to be returned = B x (1 + r) = B x (1 + 0) = B
  2. Value of N number of stock in up state = N x Su = N x 4 = 4N
  3. Total value of the portfolio = 4N - B
  4. Since this is replicating portfolio, its value must be equal to the value of the call option in the up state = Cu = max (Su - K, 0) = max (4 - 2.5, 0) = 1.5

So, our first equation is: 4N - B = 1.5

Value of the replicating portfolio at the end of 1 year in the down state:

  1. Borrowed amount to be returned = B x (1 + r) = B x (1 + 0) = B
  2. Value of N number of stock in down state = N x Sd = N x 1 = N
  3. Total value of the portfolio = N - B
  4. Since this is replicating portfolio, its value must be equal to the value of the call option in the down state = Cd = max (Sd - K, 0) = max (1 - 2.5, 0) = 0

So, our second equation is: N - B = 0

Hence, B = N

Put it in the first equation: 4N - B = 4N - N = 3N = 1.5

hence, N = 1.5 / 3 = 0.5

and B = N = 0.5

Thus, we can replicate the call option by borrowing 0.50 and buying 0.50 number of stock.

So the value of the call option today = value of the replicating portfolio today = N x S0 - B = 0.5 x 2 - 0.5 = 0.50

--------------------

Part (b): Replicating portfolio for the put option

let's create the replicating portfolio as:

  1. Lend an amount B at risk free rate, r = 0 (question states assume interest rates are zero)
  2. Short sell N number of underlying stock

Value of the replicating portfolio at the end of 1 year in the up state:

  1. Lent amount to be received back = B x (1 + r) = B x (1 + 0) = B
  2. Value of N number of stocks short sold, in up state = N x Su = N x 4 = 4N
  3. Total value of the portfolio = B - 4N
  4. Since this is replicating portfolio, its value must be equal to the value of the put option in the up state = Pu = max (K - Su, 0) = max (2.5 - 4, 0) = 0

So, our first equation is: B - 4N = 0. Hence, B = 4N

Value of the replicating portfolio at the end of 1 year in the down state:

  1. Lent amount to be received back = B x (1 + r) = B x (1 + 0) = B
  2. Value of N number of stock short sold, in down state = N x Sd = N x 1 = N
  3. Total value of the portfolio = B - N
  4. Since this is replicating portfolio, its value must be equal to the value of the call option in the down state = Pd = max (K - Sd, 0) = max (2.5 - 1, 0) = 1.50

So, our second equation is: B - N = 1.50; since B = 4N from first equation, hence B - N = 4N - N = 3N = 1.50

hence, N = 1.50 /3 = 0.5 and B = 4N = 4 x 0.5 = 2

Thus, we can replicate the put option by lending 2 and short selling 0.50 number of stock.

So the value of the put option today = value of the replicating portfolio today = B - N x S0 = 2 - 0.5 x 2 = 1.00